OpenStudy (anonymous):

Pythagorean Theorem Problem- Two planes take off at the same time from the same airport. One flies due south at 180 km/hr and the other flies due east at 240 km/hr. How far apart are they at the end of 45 minutes? Thanks!

OpenStudy (anonymous):

first figure out how far they went then draw the triangle

OpenStudy (spring98):

@Agl202

OpenStudy (anonymous):

I think the biggest issue I'm having with Pythagorean Theorem is picturing the triangle. After I see the picture I know exactly what to do. I know for sure 252 and 39 miles are two different sides of the triangle. But which sides?

OpenStudy (anonymous):

i don't get either of those numbers how did you get them?

OpenStudy (anonymous):

Sorry i meant 45 and 180 km/hr. I was looking at another problem.

OpenStudy (anonymous):

you have to figure out how far each went first

OpenStudy (anonymous):

180 km per hour for 45 minutes how far did it go?

OpenStudy (anonymous):

Well 45 minutes is 3/4 of an hour: 180 km/hr * 3/4 = 135/4 or 33.75 km. If that's what you're asking. So I would also do that for the second plane going 240 km/hr?

OpenStudy (anonymous):

yes but your arithmetic is way off

OpenStudy (anonymous):

what is \[180\times \frac{3}{4}\]?

OpenStudy (anonymous):

looks like you divided by 4 twice or something

OpenStudy (anonymous):

Christ, lol I am looking at multiple problems and by mistake put 45 kl/hr multiplied by 3/4. 180 * 3/4 is 135 kl/hr.

OpenStudy (anonymous):

Also giving that 240 * 3/4 is 180

OpenStudy (anonymous):

ok so by pythagoras there distance is \[\sqrt{135^2+180^2}\]

OpenStudy (anonymous):

*their

OpenStudy (anonymous):

Ah yes, 225. Thank you.